# Identifying Reliable Predictions in Detection Transformers

### v1

@@ 2.1 Detection Transformers (DETRs)
The encoder follows the common structure of standard computer vision models and is based on pre-trained models, whose reliability has been relatively widely explored [27, 26, 29, 20]. This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [7, 29] and distance-based out-of-distribution (OOD) detection methods [12, 28, 4]. However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction. Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of the decoder in DETR for object detection tasks.
(a) A prediction in high-reliability image: Horse
(b) A prediction in low-reliability image: Chair
Figure 2: Visualizations of the predictions generated by Cal-DETR. The optimal positive prediction (indexed by 0 and bordered in green) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented. For each prediction and layer, the maximum confidence score and its corresponding label are visualized. When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low. On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration. However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
2.2 Identifying Reliable Predictions in DETR
Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{T}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects. To address the issue, during model training, a bipartite matching algorithm is used to find the optimal matching prediction for each ground truth object based on the alignment of the class label and bounding box (as detailed in Appendix B). Consequently, the parameters $\theta$ are optimized to enhance the accuracy of these matched queries. In this paper, we refer to the matched queries as optimal positive queries, while the remaining queries are termed optimal negative queries.
At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive queries remain unknown. This raises the question of whether all predictions can be trusted and used without any post-processing, or if only a subset should be selected for the final inference—and, if so, which subset? If all of the model’s predictions, including negative queries, were generated independently and were nearly identical to one of the positive predictions, the large number of predictions would not be a concern. We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy. Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive query (i.e., prediction) per object, while assigning low scores but similar bounding boxes to negative queries (see Section 3.1 and Figure 2). As a result, the overall prediction-level calibration error becomes significantly high (See Section 3.2) if we include those noisy predictions in the final inference results. Hence, improper handling of negative queries (e.g., by using top-100 predictions) can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios, such as pedestrian detection in autonomous vehicles. For this reason, addressing this challenge and effectively distinguishing the positives from negatives is a crucial task for ensuring reliable UQ in DETR.
3 Analysis and Key Insights
3.1 Exploring the Anatomy of DETR’s Predictions
We begin by closely examining the underlying dynamics of DETR decoders and visualizing the generated outputs. Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward. We address this by reapplying the final feedforward network that operates on the last layer, $f_{\phi}$, to the intermediate layers. This allows us to transform each representation vector into its associated bounding box and class label. This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [2]. Sample visualizations are in Figure 2 along with Figures 5 and 6 in Appendix C.
In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions. In this early stage, the distinction between positive and negative queries can be ambiguous (e.g., Figure 2(a)). However, the self-attentions through the subsequent decoder layers progressively refine these predictions. By the final layer, the model selects a single query (i.e., optimal positive) and assigns a confidence score based on its understanding of the image and the object. In contrast, the confidence scores for neighboring queries (i.e., optimal negatives) do not increase to the same extent as the positives and even decrease in high-reliability images. In contrast, in low-reliability images (e.g., Figure 2(b)), the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged. Based on this observation, we present our claim:
Claim 1.
Predictions from DETR within a given image exhibit varying levels of reliability. For each object in the image, the optimal positive (i.e., best-matched) prediction is calibrated, while the remaining optimal negative predictions are uncalibrated.
3.2 Analysis of Calibration Error for Positive and Negative Predictions
Primarily, to address the first research question (RQ1) and provide quantitative support for our claim regarding the varying levels of reliability across predictions, we evaluate and compare the calibration errors on the optimal positive predictions and the optimal negative predictions. The optimal positives and negatives are determined using the standard bipartite matching algorithm [1].
We evaluate the object-level calibration error, which is defined later in Section 4.2. From Table 1, we confirm our claim that only the optimal positive predictions are properly-calibrated. Furthermore, as illustrated in Figure 2, the optimal negative predictions exhibit a significantly high calibration error, driven by consistently low confidence scores regardless of their actual accuracy. It is also noteworthy that even Cal-DETR, the state-of-the-art calibrated DETR variant, demonstrates poor calibration quality for optimal negative predictions.
Moreover, as will be explained in more detail later, Table 3 further supports our claim by illustrating that distinguishing between positive and negative predictions becomes particularly crucial when applied to image-level UQ in DETR. In short, when confidence scores are averaged without distinguishing positive queries (see the row of $\mathsf{Conf}^{+}$ using the top-$100$ predictions), the correlation between the confidence scores and image-level reliability (i.e., precision) becomes negative. This is because the average confidence scores for negative queries are shown to be inversely correlated with image-level reliability, and the majority of DETR’s outputs are actually negative. These results further highlight the importance of employing an appropriate separation scheme to accurately identify optimal positive predictions, even when ground truth annotations are unavailable.
Table 1: Object-level calibration error on top-$100$, optimal positive, and negative predictions are reported across various models and datasets. For each model, the smallest and largest errors are highlighted in blue and red, respectively. Notably, negative queries are uncalibrated. Since the top-$100$ predictions often include these negative queries, the resulting calibration error tends to be high.
Model Predictions COCO City Foggy
D-DETR Top-100 0.582 0.552 0.618
Negative 0.803 0.789 0.844
Positive 0.275 0.341 0.419
UP-DETR Top-100 0.660 0.689 0.709
Negative 0.872 0.871 0.883
Positive 0.438 0.494 0.546
Cal-DETR Top-100 0.655 0.668 0.688
Negative 0.867 0.859 0.871
Positive 0.425 0.463 0.513
DINO Top-100 0.635 0.693 0.713
Negative 0.953 0.959 0.962
Positive 0.411 0.480 0.544
4 A Systematic Framework for Identifying Positive Predictions
Identifying the positive predictions is crucial for the reliable use of DETR. Nonetheless, identifying optimal positive predictions during the inference stage is not feasible due to the lack of ground truth labels. Therefore, an alternative systematic framework is essential not only for improving its interpretability but also for ensuring reliability. Given that the optimal matching achieves the smallest calibration error, we use calibration error to measure the quality of a separation scheme, i.e., a separation scheme with a lower calibration error is deemed a better scheme.
(a) Cal-DETR
(b) UP-DETR
Figure 3: Impact of confidence threshold selection on various performance metrics in Cal-DETR and UP-DETR on COCO.
4.1 Limitations of Existing Metrics
To this end, this section analyzes the effectiveness of existing metrics, including AP, D-ECE, LA-ECE0, and LRP, for measuring calibration error. Specifically, we modify a variable (e.g., the confidence score threshold) to generate different subsets of DETR predictions. We then assess the performance of each metric for these subsets and determine the variable value that yields the highest performance. Lastly, we evaluate the extent of separation by comparing the predicted positive set to the optimal positive set. Sample analysis results are illustrated in Figure 3. For additional results and detailed information on these metrics, please refer to Appendix D.
Primarily, as noted in several studies [18, 19, 11] and our empirical findings, the optimal AP is achieved when the threshold is set to $0.0$; AP does not penalize harshly for having low confidence predictions. However, as discussed earlier, using the entire prediction set carries a high risk of including uncalibrated negatives, leading to unreliable decisions in practical applications. Furthermore, using hundreds of predictions harms the interpretability of the model.
In contrast, the optimal ECEs are often achieved when the threshold is set close to $1.0$, meaning ECEs favor retaining fewer predictions with high confidence. This is a structural pitfall that prediction-level ECEs commonly face [11]. Since ECEs do not penalize missed detections (i.e., false negatives), they can achieve near-zero error when the evaluation positive set consists solely of highly accurate and confident predictions. Therefore, unless the model is trained to be excessively overconfident—which is unlikely given that DETR is trained on large datasets using various auxiliary loss functions—ECEs can result in very small error values when using a large confidence threshold.
LRP [18], a localization-aware performance metric, can be used instead. However, LRP is not designed as a calibration metric and does not explicitly consider calibration error. As a result, the ranking scored by LRP across the different models is not necessarily aligned with the model’s calibration qualities. See Section 6.1 for the numerical analysis and more discussions.
4.2 Proposed Metric: Object-Level Calibration Error
Notation.
For each image $\bm{x}_{i}$ having $N_{i}$ objects, we consider the ground truth set $\mathcal{T}_{\bm{x}_{i}}=\{o_{i,j}=(l_{i,j},\bm{b}_{i,j})\}_{j=1}^{N_{i}bu}$, and a set of DETR predictions: $\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})=\{t_{i,q}=(\hat{\bm{p}}_{i,q},\hat{\bm{b}}_{i,q})\}_{q=1}^{M}$ where $M$ represents the number of object queries. Here, $l_{i,j}\in[1,C]$ and $\hat{\bm{p}}_{i,q}\in\{0,1\}^{C}$ represents the ground-truth label and predicted probability distribution over $C$ classes, respectively, while $\bm{b}_{i,j}$ and $\hat{\bm{b}}_{i,q}\in\{0,1\}^{4}$ corresponds to the ground truth and predicted scaled bounding box, respectively.
Definition 1.
Consider a subset of predictions, $\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\subseteq\hat{\mathcal{T}}_{\theta}(\bm{x}_{i})$, that is generated by post-processing algorithm $\mathcal{S}$ from the entire prediction set: $\hat{\mathcal{S}}_{\theta}=\mathcal{S}\circ\hat{\mathcal{T}}_{\theta}$. We define an object-level calibration error (OCE) as the average Brier score per object:
$\mathsf{OCE}\big{(}\hat{\mathcal{S}}_{\theta};\mathcal{I}\big{)}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(i,j)\in\mathcal{I}}\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}$ (1)
$\mathsf{Brier}\big{(}\hat{\mathcal{S}}_{\theta}(\bm{x}_{i});o_{i,j}\big{)}=\sum_{c=1}^{C}\big{(}\mathbbm{1}(c=l_{i,j})-\widetilde{p}_{i,j}[c]\big{)}^{2}$ (2)
$\widetilde{p}_{i,j}[c]=\frac{1}{|\mathcal{Q}_{i,j}|}\sum_{q\in\mathcal{Q}_{i,j}}\hat{\bm{p}}_{i,q}[c]$ (3)
where $\mathcal{I}=\cup_{i}\{(i,j)\}_{j=1}^{N_{i}}$ is a set of all objects indices and $\hat{\bm{p}}(\cdot)[c]$ outputs the probability of $c$-th class; $\mathcal{Q}_{i,j}$ is a set of query indices that matches to the ground truth object $o_{i,j}$ and we propose two variants:
$\mathcal{Q}_{i,j}$$\triangleq\{q\mid\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\geq\epsilon\}~{}$ $(\mathsf{OCE_{ENS\epsilon}})$ (4)
$\mathcal{Q}_{i,j}$$\triangleq\{q=\text{argmin}_{q}\mathsf{IoU}(\bm{b}_{i,j},\hat{\bm{b}}_{i,q})\}~{}$ $(\mathsf{OCE_{MAX}})$ (5)
The difference is that $\mathsf{OCE_{ENS\epsilon}}$ ensembles the overlapping predictions, while $\mathsf{OCE_{MAX}}$ selects the prediction with the best bounding-box matching. When $|\mathcal{Q}_{i,j}|=0$, we consider the predicted probability to be zero, thus the corresponding Brier score is estimated as $1.0$. Following [14, 11], we use IoU thresholds of $\epsilon=0.5,0.75$ and report the average score as $\mathsf{OCE_{ENS}}$ or simply $\mathsf{OCE}$.
The introduced calibration error has two desirable characteristics. First, the prediction set achieves the lowest calibration error when the predictions are well-calibrated to the respective closest ground truth objects. Second, it penalizes the subset, $\hat{\mathcal{S}}_{\theta}(\cdot)$, that includes missing ground truth objects; this ensures that subsets containing a small set of highly precise predictions are not assigned an artificially low error, unlike D-ECE and LA-ECE metrics. Thus, this metric can effectively assess not only whether the given subsets $\{\hat{\mathcal{S}}_{\theta}(\bm{x}_{i})\}_{i}$ are reliable, but also whether they comprehensively capture all ground truth objects, providing richer information compare to existing metrics.
5 Use-Case Implications
5.1 Image-Level Reliability
@@
As a main use-case implication of this research, this paper investigates how effective separation impacts the performance of quantifying image-level reliability within the DETR framework. Specifically, we first introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image.
@@
$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{T}}_{\theta}(\bm{x}),~{}\mathcal{T}_{\bm{x}}\big{)}.$ (6)
@@
where one can quantify $\mathsf{Perf}$ using any standard performance metrics such as average precision, depending on the specification.
@@
By its definition, image-level reliability directly addresses the model’s applicability to a given test instance. However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible. Therefore, our goal is to employ a proxy method that quantifies uncertainty or confidence, which should closely correlate with image-level reliability.
5.2 Proposed Framework: Quantifying Reliability by Contrasting
As illustrated, positive and negative queries exhibit varying levels of reliability. Interestingly, having queries with low confidence scores does not necessarily imply low reliability. Our empirical observations show that confidence scores in negative queries actually inversely correlate with image-level reliability. More specifically, for a reliable instance, we observe that the confidence of positive queries increases across the decoder layers, while that of negative queries remains low; this results in a large gap between positive and negative queries. In contrast, for unreliable instances (e.g., Figure 2(a)), the confidence score of the positive query does not increase across the layers, whereas the scores of negative queries are either slightly elevated or remain unchanged. Consequently, there is a small gap between positive and negative queries (e.g., Figure 2(b)).
$\mathsf{ContrastiveConf}(\bm{x})$ $=\mathsf{Conf}^{+}(\bm{x})-\lambda\mathsf{Conf}^{-}(\bm{x})$ (7)
$\mathsf{Conf}^{+}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})}\mathsf{c}(t)$ (8)
$\mathsf{Conf}^{-}(\bm{x})$ $=\frac{1}{|\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{t\in\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})}\mathsf{c}(t)$ (9)
where $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{T}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor, that can be selected from the validation set. We include an ablation study of the scaling factor in the supplementary material. Our results show that the proposed method is not too sensitive to the choice of scaling factor and consistently outperforms baselines, with a scaling factor of $10.0$ providing adequate performance. $\mathsf{c}(\cdot)$ denotes the confidence estimate for the prediction; in this paper, we use maximum probability, $\mathsf{c}(t=(\hat{\bm{p}},\hat{\bm{b}}))=\max_{c}\hat{\bm{p}}[c]$.
As shown so far, identifying a $\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{T}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach. In practice, however, neither the ground truth (i.e., the queries that DETR intends to use) nor the optimal positives are available during the UQ stage. Instead, we approximate the ground truth separation by applying a post-processing algorithm that minimizes the calibration error on the validation set:
$\hat{\mathcal{T}}_{\theta}^{+}(\bm{x})=\mathcal{S}^{*}\circ\hat{\mathcal{T}}_{\theta}(\bm{x})$ (10)
$\mathcal{S}^{*}=\text{argmin}_{\mathcal{S}}\mathsf{OCE}(\mathcal{S}\circ\hat{\mathcal{T}}_{\theta};\mathcal{I}_{val})$ (11)
where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{O_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$. Unless specified otherwise, we adopt confidence thresholding for the separation scheme (i.e., post-processing algorithm), where the corresponding threshold is the variable to optimize.
Table 2: Comparison of the effectiveness of the proposed OCE and baseline metrics, focusing on their ranking correlation with the rankings given by the models’ calibration quality on optimal positive predictions (denoted by the superscript *). For each metric, various post-processing schemes (i.e., ways to identify positive predictions) are applied and indicated within parentheses. Correlations exceeding $0.9$ are highlighted in blue, while negative correlations are highlighted in red. Our OCE-based rankings consistently provide strong correlations with optimal ranking and OCE effectively measures the model’s calibration quality alongside the post-processing scheme employed.
[…]
References D-ECE 1.000 0.995 0.969 1.000 0.996 0.993 1.000 0.873 0.996 0.980
[…]
$\mathsf{OCE_{ENS{}}}^{*}$ 0.969 0.948 1.000 0.993 0.990 1.000 0.996 0.836 1.000 0.970
Baselines AP (Top-100) [6, 14] 0.552 0.583 0.343 0.779 0.784 0.700 0.597 0.843 0.523 0.634
LRP (LRP) [18] 0.103 0.051 0.338 -0.268 -0.272 -0.150 -0.257 -0.610 -0.170 -0.137
D-ECE (Top-100) -0.873 -0.833 -0.966 0.608 0.576 0.511 -0.254 0.250 -0.316 -0.144
D-ECE (0.3) [10] 0.905 0.939 0.827 0.982 0.967 0.990 0.964 0.799 0.981 0.928
D-ECE (LRP) 0.802 0.855 0.659 0.987 0.984 0.999 0.983 0.793 0.996 0.895
D-ECE ($\mathsf{OCE_{ENS}}$) 0.856 0.901 0.727 0.987 0.983 0.999 0.983 0.793 0.996 0.914
LA-ECE0 (Top-100) -0.944 -0.913 -0.994 -0.431 -0.482 -0.517 -0.786 -0.386 -0.822 -0.697
LA-ECE0 (0.3) -0.789 -0.729 -0.842 -0.780 -0.818 -0.830 -0.451 -0.041 -0.468 -0.639
LA-ECE0 (LRP) [11] -0.908 -0.864 -0.948 -0.833 -0.877 -0.846 -0.901 -0.597 -0.913 -0.854
LA-ECE0 ($\mathsf{OCE_{ENS}}$) -0.901 -0.855 -0.943 -0.723 -0.780 -0.739 -0.884 -0.568 -0.897 -0.810
Ours $\mathsf{OCE_{ENS}}$ ($\mathsf{OCE_{ENS}}$) 0.962 0.940 1.000 0.987 0.983 0.999 0.964 0.737 0.984 0.951
$\mathsf{OCE_{MAX}}$ ($\mathsf{OCE_{MAX}}$) 0.954 0.929 0.998 0.954 0.958 0.983 0.955 0.697 0.976 0.934
6 Experiments

### v2

@@ III-A Detection Transformer (DETR)
It is noteworthy that the encoder follows the common structure of standard computer vision models, whose reliability has been relatively widely explored [26, 27, 28, 29]. This foundation further enables the use of prominent post hoc UQ techniques, such as Monte Carlo dropout [30]. However, despite the decoder being the predominant component for object detection, there is a gap in understanding and quantifying its reliability due to its unique structural characteristic: set prediction. Therefore, this paper delves into the underlying characteristics of these predictions and presents a methodology to quantify the reliability of DETR.
III-B Bipartite Matching
Since the number of queries in DETR, $M$, is much higher than the number of annotated objects, DETR matches each object with the corresponding best model prediction during its training. To compute this optimal matching for the predictions in a given image, a bipartite matching algorithm is applied. More specifically, a matching cost between each pair of a given prediction and an object is defined as follows:
$\mathcal{L}_{matching}=\mathcal{L}_{class}+\mathcal{L}_{box}$ (1)
where $\mathcal{L}_{class}$ is the negative prediction confidence of the ground truth class and $\mathcal{L}_{box}$ is the linear combination of the $\ell_{1}$ loss between the corners of the bounding boxes and $\mathcal{L}_{iou}$. $\mathcal{L}_{iou}$ is the Generalized Intersection over Union (GIoU) [31] loss between bounding boxes. After computing this matching cost for every combination of prediction set and ground truth objects, DETR then efficiently calculates the permutation that minimizes the total matching cost using the Hungarian matching algorithm [32, 7].
III-C Expected Calibration Error (ECE) Metrics
D-ECE. For a given object detection model, detection expected calibration error (D-ECE) [8] quantifies how closely the model’s predicted confidences align with its observed precision. Specifically, let $\hat{\mathcal{D}}$ be the set of all detections that the model produces on the validation set. Each detection has an associated predicted confidence score $\hat{p}\in[0,1]$. To compute D-ECE, the confidence space $[0,1]$ is partitioned into $J$ bins, for instance via uniform intervals $\big{[}0,\tfrac{1}{J}\big{)},\big{[}\tfrac{1}{J},\tfrac{2}{J}\big{)},\dots,\big{[}\tfrac{J-1}{J},1\big{]}$. Let $\widehat{\mathcal{D}}_{j}\subseteq\widehat{\mathcal{D}}$ be the subset of detections falling into the $j$-th bin. Denote by $\bar{p}_{j}$ the average confidence of detections in bin $j$, and by $\mathrm{precision}(j)$ the empirical precision in bin $j$. A detection is considered a true positive if it has the correct predicted class label and its IoU with the ground-truth box exceeds a given threshold $\tau$. The empirical precision $\mathrm{precision}(j)$ is thus the fraction of detections in $\widehat{\mathcal{D}}_{j}$ that are true positives.
D-ECE then aggregates the absolute difference between each bin’s average confidence and observed precision, weighted by bin size, and is defined as:
[…]
Following the standard protocol [20], two IoU thresholds ($\tau=0.5$ and $\tau=0.75$) are used, and the D-ECE scores obtained under these thresholds are averaged to yield the final metric. Because this process is carried out on a validation set, it provides an overall measure of the model’s calibration performance on that set. A lower D-ECE indicates that the confidence scores are better aligned with actual precision, while a higher D-ECE reveals a potential mismatch between confidence and true detection performance.
LaECE. To account for both localization error and classification precision, localisation-aware expected calibration error (LaECE) [19] aligns model confidence with the product of the predicted bounding box’s precision and its IoU with the ground-truth box. Specifically, LaECE is computed in a class-wise manner to mitigate class imbalance. Let $c$ index the class, and let $\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$. Partition the confidence scores into $J$ bins, and let $\widehat{\mathcal{D}}_{j}^{c}\subseteq\widehat{\mathcal{D}}^{c}$ be the set of detections for class $c$ in bin $j$. Define $\bar{p}^{c}_{j}$ to be the average confidence in $\widehat{\mathcal{D}}_{j}^{c}$, $\mathrm{precision}^{c}(j)$ to be the corresponding precision, and $\bar{\mathrm{IoU}}^{c}(j)$ to be the average IoU of $\widehat{\mathcal{D}}_{j}^{c}$. The LaECE metric is then given by:
$\text{LaECE}\triangleq\frac{1}{K}\sum_{c=1}^{K}\sum_{j=1}^{J}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}_{j}^{c}-\text{precision}^{c}(j)\times\bar{\text{IoU}}^{c}(j)\Bigr{|}$ (3)
where $K$ is the total number of classes, and an IoU threshold $\tau$ is employed to determine whether a detection is considered a true positive.
LaECE0 [20] sets the IoU threshold $\tau$ to zero and assigns an average IoU of zero to false positive detections, thereby reducing the original formulation. In particular, if a detection is not a true positive under $\tau=0$, its IoU contribution is taken to be zero. The resulting metric becomes:
$\text{LaECE}_{0}=\frac{1}{K}\sum^{K}_{c=1}\sum^{J}_{j=1}\frac{|\mathcal{\hat{D}}_{j}^{c}|}{|\mathcal{\hat{D}}^{c}|}\cdot\Bigl{|}\bar{p}^{c}_{j}-\bar{\text{IoU}}^{c}(j)\Bigr{|}.$ (4)
By setting $\tau=0$, false positives have $\bar{\mathrm{IoU}}=0$, and thus the calibration criterion focuses more explicitly on the contribution of bounding box localization to the overall calibration error.
III-D Image-Level Reliability
@@
We introduce a formal definition of image-level reliability by examining the model’s overall object detection performance on the image, extending [33, 29].
@@
$\mathsf{ImReli}(\bm{x};\theta)\triangleq\mathsf{Perf}\big{(}\hat{\mathcal{D}}_{\theta}(\bm{x}),~{}\mathcal{D}_{\bm{x}}\big{)}$ (5)
@@
where one can use any standard performance metrics, such as average precision and recall, for $\mathsf{Perf}$ depending on the user’s requirements.
@@
By its definition, image-level reliability directly addresses the model’s applicability to a given test instance. However, since image-level reliability requires ground truth annotations for its determination, obtaining it during inference is not feasible. Therefore, this paper’s objective is to develop a method that assigns a quantitative score to each image instance, closely aligning with the model’s image-level reliability.
(a) Predictions within High-Reliability Image
(b) Predictions within Low-Reliability Image
Figure 3: Visualizations of the predictions generated by Cal-DETR. The optimal positive prediction (indexed by 0 and bordered in blue) and the five optimal negative predictions (indexed by 1-5 and bordered in red) with the largest IoU are presented. For each prediction and layer, the maximum confidence score and its corresponding label are visualized. When the model is confident, the confidence score of the positive prediction either increases or remains high across the decoder layers, while those of negative predictions decrease or stay low despite its accurate bounding box predictions. On the other hand, when the model is uncertain, DETR assigns a confidence score to positive prediction based on its confidence, thereby maintaining good calibration. However, conversely, it slightly increases or maintains the confidence scores for negative predictions.
IV Impact of Negative Predictions on DETR’s Reliability
IV-A Motivation and Scope
Notably, the number of predictions generated by DETR, $|\widetilde{\mathcal{D}}_{\theta}(\bm{x})|$, is fixed and often in the hundreds, far exceeding the number of ground truth objects. To address this issue during model training, a bipartite matching algorithm is used to find the optimal matching prediction [7] for each ground truth object based on the alignment of the class label and bounding box (as detailed in Section III-B). The training loss depends on the quality of this matching, and as a result, the parameters $\theta$ are primarily optimized to enhance the accuracy of these matched predictions. In this paper, we refer to the matched predictions as optimal positive predictions, while the remaining predictions are termed optimal negative predictions.
At the inference stage in real-world scenarios, however, ground truth annotations are unavailable, meaning the optimal positive predictions remain unknown. This raises the question of whether all predictions are trustworthy, or should only a specific subset be chosen—and if so, which subset? If all of the model’s predictions were generated independently and were well-calibrated (e.g., Figure 1b), the large number of predictions would not be a concern. We could simply apply well-known algorithms like NMS to remove duplicates and resolve redundancy.
Notably, we have observed that DETR assigns well-calibrated confidence scores to only a single positive prediction per object. Meanwhile, unmatched negative predictions, which often have accurate bounding boxes, receive uncalibrated low scores (e.g., Figure 1c). As a result, the overall calibration error becomes significantly high if we include those noisy predictions in the final inference results. Hence, as detailed in the following sections, improper handling of negative predictions can severely compromise DETR’s reliability, with potentially catastrophic consequences in safety-critical scenarios. For this reason, properly distinguishing the positives from the negatives is a crucial task for ensuring the reliable use of DETR.
IV-B Exploring the Anatomy of DETR’s Predictions
We begin our analysis by visualizing and examining the outputs generated from the DETR decoders. Since the Transformer decoder outputs only representation vectors, investigating their evolution across layers is not straightforward. We address this by reapplying the final feedforward network that operates on the last layer, $f_{\phi}$, to the intermediate layers. This allows us to transform each representation vector into its associated bounding box and class label. This is feasible due to the alignment of intermediate representations, facilitated by residual connections between decoder layers [34]. Sample visualizations are in Figure 3.
In the first decoder layer, the model appears to explore the encoded image features, producing varied queries that result in various plausible predictions. In this early stage, the distinction between positive and negative queries can be ambiguous. However, the self-attentions through the subsequent decoder layers progressively refine these predictions. By the final layer, the model selects a single query (i.e., optimal positive prediction) and assigns a confidence score based on its understanding of the image and the object. In contrast, the confidence scores for neighboring queries (i.e., optimal negative predictions) do not increase to the same extent as the positives and even decrease in high-reliability images. In contrast, in low-reliability images, the confidence score of the positive query does not significantly increase, while the scores of the negative ones are either slightly raised or unchanged. Based on this observation, we present our claim:
Claim 1.
Predictions from DETR within a given image exhibit varying levels of reliability. For each object in the image, the optimal positive prediction is calibrated (i.e., reliable), while the remaining optimal negative predictions are not.
IV-C Numerical Analysis
(a) Object-level Calibration Error
(b) Correlation to Image-level Reliability
Figure 4: A visualization of the difference in calibration between positive and negative predictions on the COCO dataset. In Figure 4a, Object-Level Calibration Error (OCE) values of their confidence scores are shown, where a lower score represents better-calibrated predictions. In Figure 4b, Pearson Correlation Coefficient (PCC) values between the ground truth $\mathsf{ImReli}$ and the average confidence scores are shown, where a higher value is better.